# San Jose/Seattle, US https://bytedance.github.io/vidi-website/

### v1

@@ 4 Experiment
To validate the effectiveness of Vidi2, we evaluate it on three representative tasks: Temporal Retrieval, Spatio-Temporal Grounding, and Video Question Answering. While the first two are tested on our proposed benchmarks (VUE-TR-V2 and VUE-STG), the latter leverages public video QA datasets [wang2025lvbench, wu2024longvideobench, fu2025video] to assess general multimodal understanding.
@@ 4.1 Spatio-Temporal Grounding
We compare our Vidi2 model against leading large multimodal models (LMMs), including Gemini 3, GPT-5 and Qwen3-VL, on the VUE-STG benchmark. Since our system predicts full spatio-temporal tubes, we conduct a comprehensive analysis across both temporal and spatial dimensions. The metrics follow the definitions in Sec. 3.1.2.
@@
4.1.1 Inference Setup for Multimodal LLM Baselines
@@
We evaluate the Gemini 3 Pro (Preview) model, which accepts video input via URL. Specifically, we design a structured prompt that instructs the model to output a JSON array containing timestamps and bounding boxes, i.e.,
@@
In constructing our prompts, we follow the official Gemini API guidelines for temporal and spatial annotations: the MM:SS timestamp format is based on the Video Understanding specification  [^3] https://ai.google.dev/gemini-api/docs/video-understanding, while the bounding box coordinate system in $[0,1000]$ is defined in the Image Understanding documentation  [^4] https://ai.google.dev/gemini-api/docs/image-understanding.
@@
GPT-5 API does not accept video directly; instead, it supports a sequence of up to $120$ image frames per request. We sample frames according to the following rule: (1) for videos shorter than two minutes, we extract frames at $1$ FPS, matching Vidi2’s input configuration; (2) for longer videos, we uniformly subsample frames to satisfy the $120$-frame limit. The prompt for GPT-5 is
@@
For Qwen3-VL, we test the Qwen3-VL-32B-Instruct model following the spatio-temporal grounding examples provided in the Qwen3-VL blog [^5] https://qwen.ai/blog?id=qwen3-vl and the interleaved timestamp-image pattern demonstrated in the official cookbook [^6] https://github.com/QwenLM/Qwen3-VL/blob/main/cookbooks/video_understanding.ipynb.
@@
The four models generate temporal grounding signals in mutually incompatible representations. To enable fair comparison, we convert all predictions into a unified timeline measured in seconds. Specifically, 1) Gemini 3’s MM:SS timestamps are parsed and converted into seconds; 2) GPT-5’s frame indices are mapped to seconds according to the sampling policy; 3) Qwen3-VL directly reports timestamps in seconds, requiring no additional temporal transformation. After conversion, all timestamps are discretized by rounding to the nearest integer second to obtain a consistent per-second temporal representation.
@@ 4.1.2 Performance Evaluation
Table 5 summarizes the temporal grounding results across different video lengths, ground-truth temporal durations, and object sizes. Vidi2 establishes a clear advantage in all conditions, achieving the highest temporal precision ($73.00$), recall ($59.80$), and IoU ($53.19$). The gap is especially pronounced on medium-length videos, where Gemini 3 Pro (Preview), GPT-5 and Qwen3-VL-32B degrade significantly. These findings highlight that Vidi2 maintains robust temporal alignment even under long-duration or fine-grained temporal conditions.
@@
Table 6 reports the spatio-temporal grounding accuracy. Across all categories, Vidi2 again delivers the strongest results, with the highest vP ($44.56$), vR ($36.32$), and vIoU ($32.57$), as well as a large margin on the intersection-based metric vIoU-Int ($60.30$). These improvements indicate that Vidi2 not only identifies the correct temporal extent but also maintains coherent spatial localization throughout the entire tube. In contrast, Gemini 3 Pro (Preview), GPT-5 and Qwen3-VL-32B exhibit severe degradation on long videos or small-object scenarios, suggesting instability in spatial tracking when reasoning over extended temporal horizons.
Major Type Category Metric(%) Vidi2 Gemini 3 Pro Prev. GPT-5 Qwen3-VL-32B
Overall Overall tIoU 53.19 27.50 16.40 25.91
tP 73.00 51.91 38.29 45.29
tR 59.80 35.26 19.53 39.19
Video Length Ultra-Short (<1m<1m) tIoU 61.43 38.68 64.66 54.33
tP 85.04 58.91 78.34 76.54
tR 66.80 52.29 75.80 69.22
Short (110m1-10m) tIoU 61.48 35.39 19.73 38.97
tP 81.83 66.49 43.86 58.66
tR 68.36 43.59 24.27 57.27
Medium (1030m10-30m) tIoU 47.27 21.13 4.10 13.19
tP 65.98 43.15 23.86 30.45
tR 54.01 27.37 4.77 23.64
Tube Duration Micro-Short (<3s<3s) tIoU 47.24 25.92 16.70 24.44
tP 56.49 37.06 30.61 44.74
tR 59.78 39.66 20.95 35.75
Ultra-Short (3s10s3s\!-\!10s) tIoU 54.66 29.51 17.63 25.80
tP 73.13 51.77 38.17 46.23
tR 60.85 38.03 21.22 40.21
Short (10s60s10s\!-\!60s) tIoU 52.14 23.20 13.23 26.85
tP 79.93 58.71 41.25 43.23
tR 57.18 26.43 14.72 38.18
Object Size Small ($10\%$) tIoU 51.46 28.35 17.57 26.01
tP 70.79 52.00 40.94 46.46
tR 57.94 37.16 20.46 38.99
Medium ($10\%-30\%$) tIoU 54.51 27.23 16.66 26.85
tP 73.85 51.58 36.91 45.04
tR 61.56 34.41 20.04 40.76
Large ($30\%$) tIoU 54.97 26.06 13.49 24.32
tP 76.55 52.23 34.77 43.18
tR 61.21 32.38 16.76 37.29
Major Type Category Metric(%) Vidi2 Gemini 3 Pro Prev. GPT-5 Qwen3-VL-32B
Overall Overall vIoU 32.57 4.61 5.47 5.12
vIoU-Int. 60.30 16.59 33.64 18.47
vP 44.56 8.95 13.01 8.61
vR 36.32 5.71 6.50 7.49
Video Length Ultra-Short (<1m<1m) vIoU 36.70 7.77 20.60 11.30
vIoU-Int. 57.86 19.44 31.36 20.56
vP 50.06 12.44 25.19 16.11
vR 39.46 9.96 23.98 13.95
Short (110m1-10m) vIoU 39.62 6.36 6.86 8.63
vIoU-Int. 62.97 18.45 34.45 21.01
vP 52.41 12.36 14.97 12.37
vR 43.45 7.61 8.43 12.56
Medium (1030m10-30m) vIoU 28.18 3.05 1.44 2.02
vIoU-Int. 59.37 14.33 34.52 15.08
vP 39.46 6.49 8.47 4.75
vR 32.10 3.84 1.68 3.55
Tube Duration Micro-Short (<3s<3s) vIoU 27.96 3.93 5.60 4.31
vIoU-Int. 57.28 12.98 32.95 16.06
vP 33.25 5.85 10.52 8.10
vR 34.74 5.20 6.74 5.34
Ultra-Short (3s10s3s\!-\!10s) vIoU 34.01 4.79 5.74 4.57
vIoU-Int. 60.98 16.30 33.40 17.32
vP 45.17 8.66 12.90 8.07
vR 37.52 6.08 6.94 7.15
Short (10s60s10s\!-\!60s) vIoU 31.03 4.48 4.76 6.85
vIoU-Int. 59.83 18.43 34.27 22.45
vP 48.00 11.03 14.14 10.16
vR 34.01 5.01 5.29 9.27
Object Size Small ($10\%$) vIoU 23.31 2.33 3.66 1.49
vIoU-Int. 44.13 7.90 21.38 5.88
vP 32.08 4.36 9.06 2.73
vR 25.71 2.85 4.18 2.17
Medium ($10\%-30\%$) vIoU 39.42 5.83 7.34 7.08
vIoU-Int. 71.21 20.41 42.19 23.65
vP 52.97 10.92 15.93 11.29
vR 44.44 7.16 8.69 10.20
Large ($30\%$) vIoU 42.50 7.77 6.61 10.09
vIoU-Int. 77.22 30.22 48.08 38.02
vP 59.14 15.86 16.96 17.08
vR 47.29 9.78 8.27 15.00
Category Metric(%)       Vidi2 Gemini 3 Pro Prev.    GPT-5
Overall $\bar{\text{IoU}}$ 48.75 37.58 17.15
$\bar{\text{P}}$ 62.45 48.61 29.64
$\bar{\text{R}}$ 64.93 56.30 26.63
Ultra-Short (<1m<1m) $\bar{\text{IoU}}$ 59.09 56.72 41.08
$\bar{\text{P}}$ 69.76 68.56 63.84
$\bar{\text{R}}$ 82.48 75.22 48.19
Short (110m1-10m) $\bar{\text{IoU}}$ 49.01 49.40 26.52
$\bar{\text{P}}$ 65.38 61.83 41.64
$\bar{\text{R}}$ 66.05 74.93 38.91
Medium (1030m10-30m) $\bar{\text{IoU}}$ 51.05 40.56 11.28
$\bar{\text{P}}$ 63.42 50.48 20.51
$\bar{\text{R}}$ 67.91 63.69 21.50
Long (3060m30-60m) $\bar{\text{IoU}}$ 48.15 26.20 9.39
$\bar{\text{P}}$ 60.54 36.73 20.21
$\bar{\text{R}}$ 62.56 39.36 16.05
Ultra-Long (>60m>60m) $\bar{\text{IoU}}$ 38.65 21.19 12.49
$\bar{\text{P}}$ 54.55 32.18 21.49
$\bar{\text{R}}$ 50.98 30.11 22.27
Keyword $\bar{\text{IoU}}$ 47.73 38.41 23.17
$\bar{\text{P}}$ 63.21 52.43 41.07
$\bar{\text{R}}$ 63.72 56.00 32.59
Phrase $\bar{\text{IoU}}$ 50.11 37.84 17.02
$\bar{\text{P}}$ 63.80 48.94 29.18
$\bar{\text{R}}$ 67.07 57.40 25.88
Sentence $\bar{\text{IoU}}$ 48.17 36.73 12.92
$\bar{\text{P}}$ 60.59 45.51 22.26
$\bar{\text{R}}$ 63.75 55.46 23.04
Audio $\bar{\text{IoU}}$ 46.90 33.27 6.93
$\bar{\text{P}}$ 57.51 39.42 12.16
$\bar{\text{R}}$ 67.09 58.26 23.21
Vision $\bar{\text{IoU}}$ 51.04 39.54 20.62
$\bar{\text{P}}$ 63.12 52.16 35.78
$\bar{\text{R}}$ 66.94 54.76 26.85
Vision+Audio $\bar{\text{IoU}}$ 46.81 37.26 17.85
$\bar{\text{P}}$ 64.03 48.66 30.17
$\bar{\text{R}}$ 61.38 57.26 28.04
Table 7: Performance of different models on the VUE-TR-V2 benchmark across various evaluation attributes. $\bar{P}$ and $\bar{R}$ denote the AUC (Area Under Curve) of precision and recall, respectively, while $\bar{\text{IoU}}$ represents the AUC of intersection-over-union between predicted and ground-truth timestamp ranges, following the definition in [team2025vidi]. GPT models are evaluated via the Azure API, and Gemini models via the official Google API. The latest Gemini 3 Pro (Preview) model is tested by directly uploading full videos (no size restriction), whereas GPT-5 is limited by the Azure API’s $120$-frame cap; for videos exceeding $120$ seconds, we uniformly sample $120$ frames.
@@ 4.2 Temporal Retrieval
We compare our model with three state-of-the-art proprietary models including Gemini 3 Pro (Preview) [comanici2025gemini] and GPT-5 [OpenAI2025GPT5SystemCard]. There models are chosen for their strong performance and broad modality support, making them competitive baselines for real-world temporal retrieval.
The latest Gemini 3 Pro (Preview) naturally supports $\ast$.mp4’ as video input with maximum length over $2$ hours. Compared with previous 0325 version, the latest version can follow the temporal retrieval task very well and the result stably contains timestamps in HH:MM:SS. One successful example prompt is used as follows.
@@
Since The lastest GPT-5 API does not support audio, we extract frames at $1$ fps and feed them as input images. Following Vidi [team2025vidi], we design a simple custom prompt with in-context examples, ensuring that the output only contains frame index ranges. An example instruction is shown as
@@
We present the detailed results in terms of length, query format, and modality subsets in Table 7. Notably, Vidi2 significantly outperforms Gemini 3 Pro (Preview) and GPT-5 for overall IoU ($10\%$). For detailed analysis on different attributes, Vidi2 also excels on IoU, precision, recall on most subsets. We can conclude that a higher overall IoU of Vidi2 score ensures a better experience for automatic video editing/trimming scenarios.
LVBench [wang2025lvbench] LongVideoBench [wu2024longvideobench] VideoMME [fu2025video]
Gemini-2.5-Pro [comanici2025gemini] 78.7 84.3 -
Qwen2.5 VL-7B [qwen2.5-VL] 45.3 54.7 65.1

### v2

@@ 4 Experiment
To validate the effectiveness of Vidi2, we evaluate it on three representative tasks: Temporal Retrieval, Spatio-Temporal Grounding, and Video Question Answering. While the first two are tested on our proposed benchmarks (VUE-TR-V2 and VUE-STG), the latter leverages public video QA datasets [19, 21, 3] to assess general multimodal understanding. VUE-PLOT is added to evaluate Vidi2 and competitive models for complex plot understanding.
@@ 4.1 Spatio-Temporal Grounding
We compare our Vidi2 model against leading large multimodal models (LMMs), including Gemini 3, GPT-5 and Qwen3-VL, on the VUE-STG benchmark. Since our system predicts full spatio-temporal tubes, we conduct a comprehensive analysis across both temporal and spatial dimensions. The metrics follow the definitions in Section 3.1.2.
@@
4.1.1 Inference Setup for LMMs
@@
We evaluate the Gemini 3 Pro Preview model, which accepts video input via URL. Specifically, we design a structured prompt that instructs the model to output a JSON array containing timestamps and bounding boxes, i.e.,
@@
In constructing our prompts, we follow the official Gemini API guidelines for temporal and spatial annotations: the MM:SS timestamp format is based on the Video Understanding specification https://ai.google.dev/gemini-api/docs/video-understanding, while the bounding box coordinate system in $[0,1000]$ is defined in the Image Understanding documentation §§§https://ai.google.dev/gemini-api/docs/image-understanding.
@@
GPT-5 API does not accept video directly; instead, it supports a sequence of up to $120$ image frames per request. We sample frames according to the following rule: 1) for videos shorter than two minutes, we extract frames at $1$ FPS, matching Vidi2’s input configuration; 2) for longer videos, we uniformly subsample frames to satisfy the $120$-frame limit. The prompt for GPT-5 is
@@
Major Type Category Metric(%) Vidi2.5 Vidi2 Gemini 3 Pro Prev. GPT-5 Qwen3-VL -32B
Overall Overall tIoU 58.34 53.19 27.50 16.40 25.91
tP 72.14 73.00 51.91 38.29 45.29
tR 68.00 59.80 35.26 19.53 39.19
Video Length Ultra-Short (<1m<1m) tIoU 73.13 61.43 38.68 64.66 54.33
tP 85.99 85.04 58.91 78.34 76.54
tR 79.89 66.80 52.29 75.80 69.22
Short (110m1-10m) tIoU 65.31 61.48 35.39 19.73 38.97
tP 79.40 81.83 66.49 43.86 58.66
tR 74.04 68.36 43.59 24.27 57.27
Medium (1030m10-30m) tIoU 51.63 47.27 21.13 4.10 13.19
tP 65.48 65.98 43.15 23.86 30.45
tR 62.38 54.01 27.37 4.77 23.64
Tube Duration Micro-Short (<3s<3s) tIoU 53.40 47.24 25.92 16.70 24.44
tP 59.48 56.49 37.06 30.61 44.74
tR 68.44 59.78 39.66 20.95 35.75
Ultra-Short (3s10s3s\!-\!10s) tIoU 58.60 54.66 29.51 17.63 25.80
tP 71.01 73.13 51.77 38.17 46.23
tR 68.25 60.85 38.03 21.22 40.21
Short (10s60s10s\!-\!60s) tIoU 59.84 52.14 23.20 13.23 26.85
tP 80.49 79.93 58.71 41.25 43.23
tR 67.17 57.18 26.43 14.72 38.18
Object Size Small ($10\%$) tIoU 56.35 51.46 28.35 17.57 26.01
tP 71.05 70.79 52.00 40.94 46.46
tR 65.55 57.94 37.16 20.46 38.99
Medium ($10\%-30\%$) tIoU 60.98 54.51 27.23 16.66 26.85
tP 73.98 73.85 51.58 36.91 45.04
tR 71.15 61.56 34.41 20.04 40.76
Large ($30\%$) tIoU 58.72 54.97 26.06 13.49 24.32
tP 71.78 76.55 52.23 34.77 43.18
tR 68.62 61.21 32.38 16.76 37.29
Major Type Category Metric(%) Vidi2.5 Vidi2 Gemini 3 Pro Prev. GPT-5 Qwen3-VL -32B
Overall Overall vIoU 38.64 32.57 4.61 5.47 5.12
vIoU-Int. 64.84 60.30 16.59 33.64 18.47
vP 47.26 44.56 8.95 13.01 8.61
vR 44.71 36.32 5.71 6.50 7.49
Video Length Ultra-Short (<1m<1m) vIoU 48.25 36.70 7.77 20.60 11.30
vIoU-Int. 64.42 57.86 19.44 31.36 20.56
vP 56.17 50.06 12.44 25.19 16.11
vR 52.05 39.46 9.96 23.98 13.95
Short (110m1-10m) vIoU 45.23 39.62 6.36 6.86 8.63
vIoU-Int. 67.78 62.97 18.45 34.45 21.01
vP 54.68 52.41 12.36 14.97 12.37
vR 50.60 43.45 7.61 8.43 12.56
Medium (1030m10-30m) vIoU 33.27 28.18 3.05 1.44 2.02
vIoU-Int. 63.29 59.37 14.33 34.52 15.08
vP 41.63 39.46 6.49 8.47 4.75
vR 40.18 32.10 3.84 1.68 3.55
Tube Duration Micro-Short (<3s<3s) vIoU 33.35 27.96 3.93 5.60 4.31
vIoU-Int. 59.05 57.28 12.98 32.95 16.06
vP 36.98 33.25 5.85 10.52 8.10
vR 41.31 34.74 5.20 6.74 5.34
Ultra-Short (3s10s3s\!-\!10s) vIoU 39.22 34.01 4.79 5.74 4.57
vIoU-Int. 65.73 60.98 16.30 33.40 17.32
vP 47.10 45.17 8.66 12.90 8.07
vR 45.32 37.52 6.08 6.94 7.15
Short (10s60s10s\!-\!60s) vIoU 39.54 31.03 4.48 4.76 6.85
vIoU-Int. 64.99 59.83 18.43 34.27 22.45
vP 52.19 48.00 11.03 14.14 10.16
vR 44.67 34.01 5.01 5.29 9.27
Object Size Small ($10\%$) vIoU 29.08 23.31 2.33 3.66 1.49
vIoU-Int. 49.86 44.13 7.90 21.38 5.88
vP 36.20 32.08 4.36 9.06 2.73
vR 33.25 25.71 2.85 4.18 2.17
Medium ($10\%-30\%$) vIoU 46.08 39.42 5.83 7.34 7.08
vIoU-Int. 73.99 71.21 20.41 42.19 23.65
vP 55.33 52.97 10.92 15.93 11.29
vR 53.37 44.44 7.16 8.69 10.20
Large ($30\%$) vIoU 48.38 42.50 7.77 6.61 10.09
vIoU-Int. 82.82 77.22 30.22 48.08 38.02
vP 59.32 59.14 15.86 16.96 17.08
vR 56.74 47.29 9.78 8.27 15.00
@@
For Qwen3-VL, we test the Qwen3-VL-32B-Instruct model following the spatio-temporal grounding examples provided in the Qwen3-VL bloghttps://qwen.ai/blog?id=qwen3-vl and the interleaved timestamp-image pattern demonstrated in the official cookbookhttps://github.com/QwenLM/Qwen3-VL/blob/main/cookbooks/video_understanding.ipynb.
@@
The four models generate temporal grounding signals in mutually incompatible representations. To enable fair comparison, we convert all predictions into a unified timeline measured in seconds. Specifically, 1) Gemini 3’s MM:SS timestamps are parsed and converted into seconds; 2) GPT-5’s frame indices are mapped to seconds according to the sampling policy; 3) Qwen3-VL directly reports timestamps in seconds, which does not require additional temporal transformation. After conversion, all timestamps are discretized by rounding to the nearest integer second to obtain a consistent per-second temporal representation.
@@ 4.1.2 Performance Evaluation
Table 6 summarizes the temporal grounding results across different video lengths, ground-truth temporal durations, and object sizes. Vidi2 establishes a clear advantage in all conditions, achieving the highest temporal precision ($73.00$), recall ($59.80$), and IoU ($53.19$). The gap is especially pronounced on medium-length videos, where Gemini 3 Pro Preview, GPT-5 and Qwen3-VL-32B degrade significantly. These findings highlight that Vidi2 maintains robust temporal alignment even under long-duration or fine-grained temporal conditions.
@@
Table 7 reports the spatio-temporal grounding accuracy. Across all categories, Vidi2 again delivers the strongest results, with the highest vP ($44.56$), vR ($36.32$), and vIoU ($32.57$), as well as a large margin on the intersection-based metric vIoU-Int ($60.30$). These improvements indicate that Vidi2 not only identifies the correct temporal extent but also maintains coherent spatial localization throughout the entire tube. Vidi2.5 further achieves significant improvement (6.07%) over Vidi2 on vIoU with additional RL training. In contrast, Gemini 3 Pro Preview, GPT-5 and Qwen3-VL-32B exhibit severe degradation on long videos or small-object scenarios, suggesting instability in spatial tracking when reasoning over extended temporal horizons.
Category Metric(%) Vidi2.5 Vidi2 Gemini 3 Pro Prev. GPT-5
Overall $\bar{\text{IoU}}$ 49.62 48.75 37.58 17.15
$\bar{\text{P}}$ 59.78 62.45 48.61 29.64
$\bar{\text{R}}$ 71.09 64.93 56.30 26.63
Ultra-Short (<1m<1m) $\bar{\text{IoU}}$ 57.60 59.09 56.72 41.08
$\bar{\text{P}}$ 66.81 69.76 68.56 63.84
$\bar{\text{R}}$ 80.66 82.48 75.22 48.19
Short (110m1-10m) $\bar{\text{IoU}}$ 50.63 49.01 49.40 26.52
$\bar{\text{P}}$ 63.61 65.38 61.83 41.64
$\bar{\text{R}}$ 73.31 66.05 74.93 38.91
Medium (1030m10-30m) $\bar{\text{IoU}}$ 51.06 51.05 40.56 11.28
$\bar{\text{P}}$ 59.94 63.42 50.48 20.51
$\bar{\text{R}}$ 73.55 67.91 63.69 21.50
Long (3060m30-60m) $\bar{\text{IoU}}$ 48.54 48.15 26.20 9.39
$\bar{\text{P}}$ 56.96 60.54 36.73 20.21
$\bar{\text{R}}$ 68.90 62.56 39.36 16.05
Ultra-Long (>60m>60m) $\bar{\text{IoU}}$ 42.22 38.65 21.19 12.49
$\bar{\text{P}}$ 53.86 54.55 32.18 21.49
$\bar{\text{R}}$ 60.73 50.98 30.11 22.27
Keyword $\bar{\text{IoU}}$ 48.56 47.73 38.41 23.17
$\bar{\text{P}}$ 59.64 63.21 52.43 41.07
$\bar{\text{R}}$ 70.63 63.72 56.00 32.59
Phrase $\bar{\text{IoU}}$ 51.22 50.11 37.84 17.02
$\bar{\text{P}}$ 61.66 63.80 48.94 29.18
$\bar{\text{R}}$ 73.10 67.07 57.40 25.88
Sentence $\bar{\text{IoU}}$ 48.83 48.17 36.73 12.92
$\bar{\text{P}}$ 58.07 60.59 45.51 22.26
$\bar{\text{R}}$ 69.47 63.75 55.46 23.04
Audio $\bar{\text{IoU}}$ 45.64 46.90 33.27 6.93
$\bar{\text{P}}$ 54.26 57.51 39.42 12.16
$\bar{\text{R}}$ 71.94 67.09 58.26 23.21
Vision $\bar{\text{IoU}}$ 52.77 51.04 39.54 20.62
$\bar{\text{P}}$ 61.37 63.12 52.16 35.78
$\bar{\text{R}}$ 73.13 66.94 54.76 26.85
Vision+Audio $\bar{\text{IoU}}$ 47.65 46.81 37.26 17.85
$\bar{\text{P}}$ 60.51 64.03 48.66 30.17
$\bar{\text{R}}$ 68.13 61.38 57.26 28.04
Table 8: Performance of different models on the VUE-TR-V2 benchmark across various evaluation attributes. $\bar{P}$ and $\bar{R}$ denote the AUC (Area Under Curve) of precision and recall, respectively, while $\bar{\text{IoU}}$ represents the AUC of intersection-over-union between predicted and ground-truth timestamp ranges, following the definition in [17]. GPT models are evaluated via the Azure API, and Gemini models via the official Google API. The latest Gemini 3 Pro (Preview) model is tested by directly uploading full videos (no size restriction), whereas GPT-5 is limited by the Azure API’s $120$-frame cap; for videos exceeding $120$ seconds, we uniformly sample $120$ frames.
@@ 4.2 Temporal Retrieval
We compare our model with three state-of-the-art proprietary models including Gemini 3 Pro Preview [2] and GPT-5 [11]. There models are chosen for their strong performance and broad modality support, making them competitive baselines for real-world temporal retrieval.
The latest Gemini 3 Pro Preview naturally supports $\ast$.mp4’ as video input with maximum length over $2$ hours. Compared with previous 0325 version, the latest version can follow the temporal retrieval task very well and the result stably contains timestamps in HH:MM:SS. One successful example prompt is used as follows.
@@
Since The latest GPT-5 API does not support audio, we extract frames at $1$ fps and feed them as input images. Following Vidi [17], we design a simple custom prompt with in-context examples, ensuring that the output only contains frame index ranges. An example instruction is shown as
@@
We present the detailed results in terms of length, query format, and modality subsets in Table 8. Notably, Vidi2 significantly outperforms Gemini 3 Pro Preview and GPT-5 for overall IoU ($10\%$). For detailed analysis on different attributes, Vidi2 also excels on IoU, precision, recall on most subsets. We can conclude that a higher overall IoU of Vidi2 score ensures a better experience for automatic video editing/trimming scenarios.
We also present the ROC curve for different metrics in Figure 7, and we can observe that Vidi2 consistently outperforms all competitors in IOU, precision and recall across the entire range of thresholds. It shows Vidi2’s strength in fine-grained temporal understanding, a critical capability for real-world video editing tasks. Vidi2.5 is slightly improved over Vidi2 on the overall IoU with RL training.
4.3 Generic Video QA
As a video understanding foundation model, Vidi2 also supports video question answering, which is a major improvement over the previous Vidi model. We evaluate Vidi2 and several popular multimodal models on three representative video QA benchmarks: LVBench [19], LongVideoBench [21], and VideoMME [3]. While Vidi2 still lags behind the leading proprietary model, Gemini-2.5-Pro [2], in general video QA performance, it achieves competitive results compared to the most popular open-source multimodal model of similar scale, namely Qwen2.5-VL-7B [16]. These results demonstrate that Vidi2 possesses solid multimodal reasoning ability despite being primarily optimized for retrieval and grounding tasks.
LVBench [19] LongVideoBench [21] VideoMME [3]
Gemini-2.5-Pro [2] 78.7 84.3 -
Qwen2.5-VL-7B [16] 45.3 54.7 65.1